### v1

# Defining A New Cross Reality: Digital Twins and Mixed Reality Worlds

Abstract
Technological advances are redefining the relationship between physical and virtual space. Traditionally, when users engage in virtual reality (VR), they are completely cut off from the physical space; similarly, they are unable to access virtual experiences while engaged in physical activities. However, modern multi-platform metaverse environments allow simultaneous participation through mobile devices, creating new opportunities for integrated experiences. This study introduces the concept of "cross-reality lifestyles" to examine how users actively combine their physical and virtual activities. We identify three patterns of integration: 1) amplification: one space enhances experiences in the other; 2) complementary: spaces offer different but equally valuable alternatives; and 3) emergence: simultaneous engagement creates entirely new experiences. By analyzing commercial platforms, we create a technical framework that addresses content design, platform infrastructure, and device interfaces. This framework guides the development of cross-reality applications while demonstrating how metaverse technologies blur the traditional boundaries between physical and virtual experiences.
journal: Publication Name
[…]
The relationship between physical and virtual space is being redefined by advances in information technology. Since the era of video games and early social virtual worlds (such as Second Life), real and virtual space have been treated as mutually exclusive: users in virtual spaces were cut off from the information of physical space, while virtual experiences were disconnected from everyday life. In recent years, advancements in immersive virtual reality technology, offering a sense of presence and enabling embodied avatar experiences, have further enhanced the ways users interact with virtual environments [1]. VR creates a sense of being in a completely different place, enabling the simulation of extraordinary, non-everyday experiences. This separation shaped the technological development in two different directions. Digital twin technology [2] positioned virtual spaces as tools for simulating and optimizing real environments, establishing physical space as primary. In contrast, metaverse platforms [3] such as VRChat [^1] https://hello.vrchat.com/ and Cluster [^2] https://cluster.mu/en created virtual worlds independent of physical constraints, creating new venues for self-expression free [4]. These platforms prioritize virtual experiences while minimizing physical connections.
However, current trends suggest a shift beyond this dualistic division between real and virtual spaces. For example, younger generations routinely combine physical activities with virtual social connections, such as doing chores while staying connected through video calls [5]. Co-existence of activities in real and virtual spaces has expanded through metaverse development, especially as platforms adopt cross-device accessibility from mobile devices, such as VR-HMD and smartphones, and IoT device integration. For example, Roblox [^3] https://www.roblox.com/, one of the world’s largest non-immersive metaverse platforms added VR-HMD support in 2023, while VRChat, which is played by many VRHMD users,introduced smartphone access in 2024. As activities in virtual spaces have grown in popularity and diverse access routes to these spaces have been developed, users who simultaneously engage in both real and virtual spaces have emerged. As seen in Fig. 1, users can participate in metaverse events while doing something in real spaces [^4] Cluster Music Jam Vol.1. https://cluster.mu/e/96cf9bff-5cd5-406b-ac50-dc51e9ef08a0. While previous studies have primarily focused on the master-slave relationship that exists between the two spaces, such as real-world VR training [6] or user activity in metaverse platforms [4], these emerging patterns of simultaneous engagement remain understudied.
Figure 1: A live music event on Cluster, a metaverse platform accessible from smartphones. Although some users are on their smartphones, they are dancing and responding to the music in VR space.
In this paper, we introduce the concept of “cross-reality life experiences” to describe emerging lifestyle patterns that arise from the seamless fusion of physical space and the metaverse. We define these experiences as novel forms of activity where neither space takes precedence over the other, and life in both realms integrates into a unified activity. These experiences encompass everything from enhancing real-world capabilities through VR to crafting personalized entertainment by harmonizing the interplay between physical and virtual actions.
We categorize these cross-reality life experiences into three integration patterns: 1) amplification, 2) complementary, and 3) emergence. Using examples from virtual platforms and communities, we analyze their distinct characteristics. Additionally, by mapping representative commercial metaverse platforms to technical requirements and integration patterns, we propose a framework for designing effective cross-reality experiences. This research offers practical insights for developers and designers aiming to create impactful and meaningful cross-reality applications.
CLASSIFICATION OF CROSS-REALITY LIFE EXPERIENCES
Milgram et al.’s Reality-Virtuality Continuum [7] positioned real and virtual spaces as poles on a spectrum, with Augmented Reality (AR), which augments physical space with virtual information, and Augmented Virtuality (AV), which augments virtual space with physical elements, filling the space between these poles. While this framework focused primarily on perceptual experiences, our research extends it to everyday activities. This extension emphasizes active user agency in combining physical and metaverse spaces.Users dynamically adjust their level of engagement based on situation and intent, moving fluidly between spaces. This concept of "cross-reality life experience" extends Mixed Reality into everyday life.
Figure 2: Patterns of integration between physical reality (R) and metaverse (M) spaces. The diagram illustrates three distinct patterns: (1) Amplification, where one space catalyzes value enhancement in the other through both reality-first and metaverse-first approaches; (2) Complementary, where M provides alternative options to overcome R’s inherent limitations while maintaining R’s core strengths; and (3) Emergence, where parallel engagement in both spaces creates qualitatively new experiences beyond simple combination.
We analyze these emerging lifestyle patterns through two dimensions: "temporal relationship," which examines whether one space takes precedence or if they coexist simultaneously, and "value interaction pattern," which explores how activities across spaces generate new forms of value. Based on these dimensions, we identify three types of integration: amplification, complementary, and emergence (Figure 2). The following sections examine each type with relevant literature.
Amplification-type Integration
Amplification-type integration involves enhancing the experiential value of either metaverse (M) or real space (R) through the intervention of the other. This pattern follows a clear temporal sequence, manifesting as either Reality-First or Metaverse-First Amplification.
Reality-First Amplification
Reality-first amplification builds upon experiences in real space, where metaverse elements enhance the inherent value. This pattern is characterized by how partial participation in the metaverse serves as a catalyst to qualitatively enhance real-world experiences. This reality-first amplification has traditionally been considered in the context of digital twins and VR training, and Slater and Sanchez-Vives provide a comprehensive overview [6].
This pattern has shown significant effects across multiple domains. Medical training, particularly surgical education, uses virtual spaces to develop complex skills in reality while ensuring safety [8, 9, 10]. VR/AR applications have also been extensively studied in tourism [11]. Importantly, virtual travel experiences complement rather than replace physical travel, increasing motivation with VR travel or enhancing real visits with AR.
In physical activities such as rehabilitation and sports, VR enables graded skill development through calibrated challenges [12]. Virtual environments slightly above current skill levels promote successful experiences that enhance real-world performance [13]. These virtual achievements build self-efficacy, which translates into improved physical outcomes.
Researchers in psychology and neuroscience has revealed new insights into real-body ownership and perception by using VR as a tool [14, 15, 16]. Research also demonstrates that embodying avatars of different races can reduce implicit bias and develop empathy [17, 18].
Metaverse-First Amplification
Metaverse-first amplification occurs when physical activities deepen the value originally created in virtual spaces. Classic examples include "anime pilgrimages," where fans visit real-world locations featured in anime or games to deepen their connection to the fictional narratives [19]. Similarly, offline "meet-ups" of virtual communities, such as online gaming groups and social networking sites, strengthen virtual bonds through physical interaction [20].
More contemporary examples include fan activities surrounding virtual YouTubers (VTubers) [21], where physical events and merchandise enhance virtual engagement. Metaverse platforms also leverage this pattern through events, connecting user-generated content (UGC) in virtual space with physical influence via streamer commentary [^5] Cluster Game Revolution Eve, 2024. https://geekaku2024.cluster.mu/, fostering broader metaverse community engagement and creative participation.
Complementary-type Integration
Complementary integration restores experiential wholeness by combining physical and metaverse options. This pattern features "simultaneously selectable states" where both spaces exist as equal alternatives, allowing users to choose optimal combinations for their needs.
Complement via Teleexistence
Telepresence robots in customer service exemplify this integration by reducing spatial and temporal constraints [22]. Virtual platforms, such as teleconferencing tools, complement face-to-face interactions by providing alternative modes of engagement. Additionlay, "silent study sessions" on Zoom provide ambient presence without social pressure [23], while screen sharing capabilities can enhance productivity beyond traditional meetings.
The COVID-19 pandemic has accelerated the adoption of virtual platforms such as Gather and Spatial Chat [24] for education and conferencing. These lightweight, browser-based 2D virtual environments, even without complex 3D capabilities, effectively complement physical spaces with unified spatial experiences.
Complement via Avatars
Avatars enable forms of self-expression that are impossible in physical reality. VTubers establish a new form of expression that goes beyond the appeal of reality by combining the inner self of a real person with a new attribute expressed through an avatar [25] Similarly, CGI-based virtual influencers are emerging as alternatives to traditional human influencers in marketing [26, 27].
On the other hand, Hatada et al. introduced a avatar system in a cafe where individuals with disabilities provide customer service using telepresence robots [28]. The robots’ standardized appearance made it difficult for employees to express their individuality, so the system incorporated customizable virtual avatars as an additional option for customer interactions. This enabled employees to seamlessly switch between physical telepresence robots and virtual avatars, allowing them to provide service from both the real and virtual spaces. The system empowered employees to engage more dynamically and authentically, creating a richer and more fulfilling customer service experience.
Emergence-type Integration
Emergent integration occurs when activities in the metaverse and physical spaces take place simultaneously and interact. Unlike complementary integration, which provides various options to users by switching across the metaverse and physical spaces, emergent integration involves the persistent coexistence and active combination of both spaces. This simultaneous integration creates a unified, multifaceted experience that would be impossible to achieve in either space alone.
Two technological advances enable this integration: the maturation of social VR communities and the widespread availability of hands-free VR devices, including VR-HMDs and smartphones. These developments make simultaneous engagement in both spaces technically feasible for general users.
Emergence with Incomplete Social Presence
The HCI and CSCW community has identified "incomplete" social presence as a significant pattern in which users modulate their level of digital engagement. Research shows that teenagers actively value this partial social presence, such as multitasking while participating video calls [5]. "Study with me" videos on YouTube Live demonstrate how users create personalized digital social environments that balance engagement with controllable peer pressure [29].
By combining "incomplete" social presence with non-immersive metaverse experiences via mobile devices, the "multitasking metaverse" experience emerges. As shown in the introduction section, users achieve optimal engagement by placing avatars in virtual spaces while maintaining physical activities. For example, in a virtual music event in the metaverse platform, users study and listen to the music while their avatar automatically dances and actively reacts at key points in the event (Fig.1). This approach satisfies users who want flexible participation without constant full immersion.
Emergence of Social Bonds
Maloney and Freeman’s report of “VR sleep”—a phenomenon where individuals sleep alone physically while maintaining a sense of social presence in a virtual environment—illustrates how the fusion of personal physical activities with virtual social interaction creates a unique form of intimacy [30]. They note that users often feel as though they are sleeping alongside friends in the virtual world, with moments blurring the boundaries between physical and virtual spaces, resulting in a hybrid experience.
Future of Emergence-type Integration
Advances in cognitive psychology with VR have expanded our understanding of embodiment and broadened the possibilities for emergent integration. AR avatar [31], which overlays virtual avatars on physical bodies through AR displays, brings the Proteus Effect [32], the observation that avatar characteristics influence users’ behavior and self-perception, into physical reality. This integration creates qualitative changes in cognitive and behavioral patterns that go beyond mere visual changes.
Research into the control of multiple virtual bodies [33] suggests possibilities beyond simple metaverse-reality combinations (M and R), introducing simultaneous engagement in multiple virtual spaces while maintaining physical presence (M1, M2, … and R).
TECHNICAL FRAMEWORK FOR CROSS-REALITY LIFESTYLE
This section presents an analytical framework for implementing cross-reality integration in metaverse environments. We first examine fundamental technical requirements across key dimensions, then analyze current platform implementations, and explore how UGC and multi-platform capabilities enable emergence-type integration.
Technical Requirements and Architecture
We establish three fundamental dimensions that characterize metaverse environments and organize their technical requirements:
  • Contents: Experience design methodology and interaction in metaverse spaces, including UGC creation and sharing systems, integration of immersive social events, and flexible participation models that adapt to user engagement levels.
  • Platform: Core technical infrastructure including multi-device support, real-time communication systems, scalability for multiple users, unified 2D/3D space management, and persistent data storage.
  • Device & Interface: Technologies bridging physical and virtual spaces through sensing, motion capture, VR/AR devices, and physical feedback systems.
Table 1 systematically organizes the relationship between these dimensions and the integration types by categorizing the technical requirement from examples shown in the previous section. For instance, while amplification-type integration requires high-precision training simulations for specific physical tasks, emergence-type integration demands functionalities supporting loose synchronous interaction between both spaces such as automated avatar control.
Table 1: Technical Requirements Matrix for Cross-Reality Integration: Mapping of Technical Dimensions against Integration Patterns.
Dimension Amplification Complementary Emergence
Contents Reality-first: [ ] High-precision training simulations for specific physical tasks (e.g., medical procedures, sports) Structured learning content that maps directly to real-world actions Context-aware guidance systems that enhance physical performance Metaverse-first: [ ] Virtual community or event (e.g., VTuber concerts) UGC management and promotion Teleconference: [ ] Meetings, seminars with stable UIs (e.g., screen sharing, easy-to-use controls) Tele-existence: [ ] Remote robot interaction or field monitoring with interactive control Avatar: [ ] Flexible avatar representation Privacy-conscious social interaction tools (e.g., Voice changer) [ ] A system without excessive manipulation in VR (e.g. automation of avatar movements and reactions) Flexible UGC tools enabling spontaneous cross-reality activities VR icon indicating physical status
Platform Reality-first: [ ] High-performance physics simulation engines Precise spatial tracking and mapping Metaverse-first: [ ] Scalable content distribution systems Cross-platform compatibility (VR, PC, mobile) Robust UGC management [ ] Low-latency, stable connections Reliable multi-user session management Compatibility across 2D/2.5D/3D environments via universal API Scaling for varying group sizes Protocols for device integration [ ] Cross-platform compatibility Always-on infrastructure supporting partial presence Dynamic resource allocation based on participation status Context-aware state synchronization
Device &
Interface
Reality-first: [ ] High-precision tracking devices Specialized input/output hardware and its integration (sensors and actuators) Metaverse-first: [ ] Consumer-grade VR/AR devices Accessible mobile interfaces Teleconference: [ ] Standard video/audio equipment Tele-existence: [ ] Remote operation control devices Environmental sensing Avatar: [ ] Motion capture for expressive virtual avatars (VTuber or streamer use) Head/face tracking or lip-sync for emotion expression [ ] Lightweight, comfortable devices for extended use Switching between Reality and Virtuality (e.g. video see-through) Multi-avatar or multi-world participation (e.g. controllers for switching input among instances) Integrated monitoring systems that track physical condition
Analysis of Commercial Platforms
We classify current metaverse platforms into six categories based on core functions and technical features, examining how each supports different forms of cross-reality integration. Note that while platforms may share features across multiple categories, we classify them according to their primary use cases and defining characteristics.
1. Industrial Metaverse Platforms (e.g. Task-specific Applications, NVIDIA Omniverse)
These platforms focus on workplace simulation and digital twin creation for the manufacturing and construction industries. They integrate CAD data and real-time sensor information to optimize plant layouts and verify operational processes. Technical requirements include precise industrial data compatibility, secure high-performance infrastructure, and precise control environments through AR/VR devices. Content priorities include accurate physical simulation with real-time capabilities, while the platform must provide enterprise-class reliability and scalability. These platforms primarily support reality-first amplification through sensor and actuator integration, and complementary integration through remote robot operation.
2. Streaming-focused Platforms (e.g., YouTube Live, Twitch)
These platforms emphasize video streaming and community building around content creators. They require high-quality video delivery and viewer engagement features (chat, donations) supported by a robust infrastructure for large-scale, simultaneous broadcasting. Technical considerations include low-latency, high-quality video delivery, real-time interaction processing, and large-scale viewer data management. Device requirements remain moderate, supporting standard PC and mobile access. These platforms effectively amplify metaverse content to external audiences. Although some streamer broadcasts contents related to emergence integration, e.g. "study with me" streaming, the primary interaction pattern remains one-way from streamer to viewer.
3. 2D Teleconference Tools (e.g., Zoom, Teams, Meet)
These communication platforms serve the needs of business meetings and online education. They prioritize stable audio and video transmission while offering document sharing and chat. Technical requirements focus on low latency, high security, intuitive user interfaces, and compatibility with standard webcams and microphones. Enterprise applications place a high priority on security and stability. These platforms represent the most mature sector of complementary integration, having established themselves as alternatives to physical meetings and becoming essential for hybrid work and online education.
4. 2D Spatial Teleconference Platforms (e.g., Gather, SpatialChat)
These platforms feature avatar movement on 2D maps with proximity-based audio functionality, enabling casual spatial interactions such as poster sessions at academic conferences and social gatherings. Technical features include 2D avatar control, distance-based audio adjustment, and lightweight browser-based implementation. By minimizing hardware requirements, these platforms enable broad user participation. As examples of complementary integration, they enhance traditional meeting systems with game-like and avatar elements to promote natural interaction. Continuous use, such as during parallel tasks, can lead to quasi-emergent integrations.
5. 3D game-based platforms (e.g. Fortnite, Roblox, Minecraft)
Originally developed as gaming platforms with predefined goals, these systems have evolved towards metaverse functionality through enhanced UGC capabilities and large-scale events. Technical requirements include advanced 3D graphics processing, support for large concurrent connections, and optimization across multiple devices (PC, game consoles, mobile). Many platforms offer proprietary development environments and marketplaces for UGC. In terms of cross-reality integration, these platforms demonstrate strong metaverse-first amplification through large-scale virtual events. Social interactions within games increasingly extend to physical activities (offline meetings, merchandise sales). While emergent community development remains limited, increasing UGC capabilities may lead to spontaneous use cases.
6. Social VR Platforms (e.g. VRChat, Cluster, SecondLife, VirtualCast, REALITY, ZEPETO)
These platforms focus on communication and world editing through 3D avatars, supporting diverse applications such as music performances, academic conferences, and community events. Technical requirements include high quality 3D avatar rendering, multi-platform support (VR/PC/mobile), and advanced body tracking technologies such as motion capture. They offer significant UGC freedom and encourage active user creation of avatars and worlds. These platforms demonstrate the diverse cross-reality integration patterns, including metaverse-first amplification through VR events extending to physical activities, complementary integration through avatar-based customer service and meetings, and emergent integration exemplified by unique cultural phenomena like VR sleep.
UGC and Multi-platform Development
Analysis shows that UGC freedom and multi-platform support play a critical role in deepening cross-reality experiences. Platforms with greater UGC freedom enable greater user creativity, leading to unexpected emergent activity. Multi-device support reduces the barriers to participation, thereby facilitating the formation of emergent communities. However, mobile platforms often face UGC limitations due to device constraints, presenting a challenge for cross-reality experience development. Future development will need to maintain experience consistency across devices, while allowing for graded UGC deployment that leverages the characteristics of each platform.
CONCLUSION
This paper introduced the concept of cross-reality lifestyles and presented a framework for analyzing the integration of physical and virtual spaces in metaverse environments. The analysis reveals three distinct patterns of integration: amplification, complementary, and emergence, and identifies the technical requirements in the dimensions of content, platform infrastructure, and device interface. A systematic analysis of commercial metaverse platforms reveals how different platform types implement these requirements and support different forms of integration.
In the future, the boundaries between platforms are expected to become increasingly fluid and new hybrid forms will emerge, such as the fusion of industrial and social functions and the integration of gaming elements and practical applications. Future research should focus on developing metrics for evaluating cross-reality integration, investigating social and psychological effects, and exploring new interaction paradigms. We hope that the framework of this research will provide guidelines for platform selection and design of technical specifications for future metaverse services and application development.
ACKNOWLEDGMENTS
This work was partially supported by JST ASPIRE Grant Number JPMJAP2327, Japan.